OpenStudy (anonymous):

Solve the given equation and indicate the number of values of θ sach that 0≤θ<2π satisfy the iquation sin(2θ)+sin(θ)=0

OpenStudy (vijay):

sin(2θ)+sin(θ)=0 2sin(θ)cos(θ)+sin(θ)=0 sin(θ)[2cos(θ)+1)=0 so, sin(θ)=0 or 2cos(θ)+1=0 ===>>>> cos(θ)=(-1/2) (θ) = 0 or (2pi/3)